2/3x - 5/6y
I got 12y/6y - 15/6y= 12-5/6y= 2-5/y
Is that correct?

Answer by Alan3354(69443)

2/3x - 5/6y
I got 12y/6y - 15/6y= 12-5/6y= 2-5/y
=======================
The DEN has to have an x & a y.
2/3x - 5/6y
Find a common denominator. The product always works
= 18xy
--------------
--> 2*6y/DEN - 5*3x/DEN
= (12y - 15x)/18xy
= (4y - 5x)/6xy